# Capsule Network Projectors are Equivariant and Invariant Learners

### v1

@@ Abstract
Learning invariant representations has been the longstanding approach to self-supervised learning. However, recently progress has been made in preserving equivariant properties in representations, yet do so with highly prescribed architectures. In this work, we propose an invariant-equivariant self-supervised architecture that employs Capsule Networks (CapsNets) which have been shown to capture equivariance with respect to novel viewpoints. We demonstrate that the use of CapsNets in equivariant self-supervised architectures achieves improved downstream performance on equivariant tasks with higher efficiency and fewer network parameters. To accommodate the architectural changes of CapsNets, we introduce a new objective function based on entropy minimisation. This approach which we name CapsIE (Capsule Invariant Equivariant Network) achieves state-of-the-art performance across all invariant and equivariant downstream tasks on the 3DIEBench dataset, while outperforming supervised baselines. Our results demonstrate the ability of CapsNets to learn complex and generalised representations for large-scale, multi-task datasets compared to previous CapsNet benchmarks. Code is available at https://github.com/AberdeenML/CapsIE.
@@ 2 Problem Statement
First to define equivariance we begin by defining a Group consisting of a set $G$ and a binary operation $\cdot$ on G, $\cdot:G\times G\rightarrow G$ such that $\cdot$ are associative; there is an identity $e$ which satisfies $e\cdot a=a=a\cdot e,\forall a\in G$; and for each $a\in G$ there exists an inverse $a^{-1}$ such that $a\cdot a^{-1}=e=a^{-1}\cdot a$. Group actions are concerned with how groups manipulate sets, where the left group action can be defined as a function $\alpha$ of group $G$ and set $S$, $\alpha:G\times S\rightarrow S$ such that $\alpha(e,s)=s,\forall s\in S$, and $\alpha(g,\alpha(h,s))=\alpha(gh,s),\forall s\in S,\text{and }\forall g,h\in G$. In our setting, we are concerned with group representations which are linear group actions acting on vector space $V$, which we define as $\rho:G\rightarrow GL(V)$ where $GL(V)$ is the general linear group on $V$. Here $\rho(g)$ describes the transformation applied to both the input data $x$ and latent $x^{\prime}$ given parameters $g$ [26, 14].
@@
As defined in [14] the goal (visually depicted in Figure 2) is to learn $f$ and $\rho_{Y}$ to construct representations that are equivariant to viewpoint transformations when $\rho_{X}$ is not known, but the group elements $g$ that parameterise the transformations are. Further details of these transformations and the benchmark 3DIEBench dataset are given in Section A.
@@ 3.1 Architecture
To align with the above problem statement and Equation 1, we aim to simultaneously learn invariant $f$ and equivariant representations by optimising our network with respect to the output activations and poses. In this case, we consider the activation vector to capture existence of semantic concepts/objects of the input, thus, the invariant information is preserved by the transformation. The pose on the other hand is designed to encode positional information related to each corresponding capsule [27] (i.e. semantic concept), therefore it contains equivariant information that that was changed by the transformation. Akin to the SIE method, we therefore, consider two embedding vectors for each image view, $z_{\text{act}}$ and $z_{\text{pose}}$, which correspond to the capsule activation vector and pose, and invariant and equivariant components respectively.
To enforce our network to learn equivariant properties we utilise a prediction network $p_{\psi,g}$ which takes as input the transformation $g$ and $z_{\text{pose}}$ to predict $z^{\prime}_{\text{pose}}$, and hence learn $\rho_{Y}(g)$. In our setting, $g\in\mathbb{R}^{3}$ corresponding to the quaternions of the rotation applied. In this work, we employ the hypernetwork approach taken by [14] which uses an MLP that takes as input the transformation parameters $g$ to parameterise an MLP predictor. Such a network avoids the case where the transformation parameters $g$ are ignored and the predictor provides invariant solutions. We present more details of the predictor network in the appendix, and visually depict the full CapsIE architecture in Figure 1.
@@ 3.3 Objective Functions
As previously stated in Section 2, our goal is to learn the predictor $p_{\psi,g}$ to model $\rho_{Y}(g)$ as to enforce equivariant representations. This is achieved by maximising the cosine similarity between the output vector of the predictor $p_{\psi,g}(Z_{\text{pose}})$ given translation parameters $g$ and equivariant representation $Z_{\text{pose}}$, and the augmented view’s equivariant representation vector $Z^{\prime}_{\text{pose}}$. To avoid collapse and improve training stability we also regularise the output of $p_{\psi,g}(Z_{\text{pose}})$ by minimising the variance of the predicted equivariance representation to avoid collapse. Whereas, SIE [14] finds this to be an optional but recommended component, we found in practice, without such regularisation the predictor would consistently collapse to trivial solutions.
@@ 4 Experimentation
Capsule Projector - 1st Intermediate Embedding
@@ 7 Conclusion
Our proposed method demonstrates how CapsNets can be employed in self-supervised learning to better learn equivariant properties without the need to explicitly split representation vectors and train separate projector networks. The resulting solution, CapsIE, achieves state-of-the-art performance in equivariant and invariant downstream benchmarks with a significant improvement of 0.05 $R^{2}$ on prior self-supervised rotation prediction tasks and 0.2 $R^{2}$ improvement over the supervised baseline. In addition, we observe competitive performance of CapsIE on equivariant tasks not explicitly trained for, further demonstrating the implicit equivariant properties of our capsule architecture under standard invariant optimisation criteria akin to those of VICReg [3], SimCLR [7], and MSN [1]. Our results contribute significantly to the avocation of CapsNets in self-supervised representation learning, introducing desirable properties with improved effectiveness over MLP projectors.
@@ B.2 Downstream Evaluation
B.3 Supervised Training of SR-Caps

### v2

@@ Abstract
Learning invariant representations has been the longstanding approach to self-supervised learning. However, recently progress has been made in preserving equivariant properties in representations, yet do so with highly prescribed architectures. In this work, we propose an invariant-equivariant self-supervised architecture that employs Capsule Networks (CapsNets) which have been shown to capture equivariance with respect to novel viewpoints. We demonstrate that the use of CapsNets in equivariant self-supervised architectures achieves improved downstream performance on equivariant tasks with higher efficiency and fewer network parameters. To accommodate the architectural changes of CapsNets, we introduce a new objective function based on entropy minimisation. This approach which we name CapsIE (Capsule Invariant Equivariant Network) achieves state-of-the-art performance across invariant and equivariant tasks on the 3DIEBench dataset compared to prior equivariant SSL methods, while outperforming supervised baselines. Our results demonstrate the ability of CapsNets to learn complex and generalised representations for large-scale, multi-task datasets compared to previous CapsNet benchmarks. Code is available at https://github.com/AberdeenML/CapsIE.
@@ 2 Problem Statement
First to define equivariance we begin by defining a Group consisting of a set $G$ and a binary operation $\cdot$ on $G$, $\cdot:G\times G\rightarrow G$ such that $\cdot$ are associative; there is an identity $e$ which satisfies $e\cdot a=a=a\cdot e,\forall a\in G$; and for each $a\in G$ there exists an inverse $a^{-1}$ such that $a\cdot a^{-1}=e=a^{-1}\cdot a$. Group actions are concerned with how groups manipulate sets, where the left group action can be defined as a function $\alpha$ of group $G$ and set $S$, $\alpha:G\times S\rightarrow S$ such that $\alpha(e,s)=s,\forall s\in S$, and $\alpha(g,\alpha(h,s))=\alpha(gh,s),\forall s\in S,\text{and }\forall g,h\in G$. In our setting, we are concerned with group representations which are linear group actions acting on vector space $V$, which we define as $\rho:G\rightarrow GL(V)$ where $GL(V)$ is the general linear group on $V$. Here $\rho(g)$ describes the transformation applied to both the input data $x$ and latent $f(x)$ given parameters $g$ [26, 14].
@@
As defined in [14] the goal (visually depicted in Figure 2) is to learn $f$ and $\rho_{Y}$ to construct representations that are equivariant to viewpoint transformations when $\rho_{X}$ is not known, but the group elements $g$ that parameterise the transformations are known. Further details of these transformations and the benchmark 3DIEBench dataset are given in Section A.
@@ 3.1 Architecture
To align with the above problem statement and Equation 1, we aim to simultaneously learn invariant and equivariant representations by optimising our network $f$ with respect to the output activations and poses. In this case, we consider the activation vector to capture existence of semantic concepts/objects of the input, thus, the invariant information is preserved by the transformation. The pose on the other hand is designed to encode positional information related to each corresponding capsule [27] (i.e. semantic concept), therefore it contains equivariant information that that was changed by the transformation. Akin to the SIE method, we therefore, consider two embedding vectors for each image view, $z_{\text{act}}$ and $z_{\text{pose}}$, which correspond to the capsule activation vector and pose, and invariant and equivariant components respectively.
To enforce our network to learn equivariant properties we utilise a prediction network $p_{\psi,g}$ which takes as input the transformation $g$ and $z_{\text{pose}}$ to predict $z^{\prime}_{\text{pose}}$, and hence learn $\rho_{Y}(g)$. In our setting, $g\in\mathbb{R}^{3}$ corresponding to the quaternions of the rotation applied. In this work, we employ the hypernetwork approach taken by [14] which uses an linear projector that takes as input the transformation parameters $g$ to parameterise an MLP predictor. Such a network avoids the case where the transformation parameters $g$ are ignored and the predictor provides invariant solutions. We present more details of the predictor network in the appendix, and visually depict the full CapsIE architecture in Figure 1.
@@ 3.3 Objective Functions
As previously stated in Section 2, our goal is to learn the predictor $p_{\psi,g}$ to model $\rho_{Y}(g)$ as to enforce equivariant representations. This is achieved by maximising the cosine similarity between the output vector of the predictor $p_{\psi,g}(Z_{\text{pose}})$ given translation parameters $g$ and equivariant representation $Z_{\text{pose}}$, and the augmented view’s equivariant representation vector $Z^{\prime}_{\text{pose}}$. To avoid collapse and improve training stability we also regularise the output of $p_{\psi,g}(Z_{\text{pose}})$ by ensuring the variance of the predicted equivariance representation is 1 to avoid collapse. Whereas, SIE [14] finds this to be an optional but recommended component, we found in practice, without such regularisation the predictor would consistently collapse to trivial solutions.
@@ 4.1 Training Protocol
Projector - 1st Intermediate Embedding
SIE 20.1M 13.07G 512 512 - 80.53 77.64
@@ 4.2 Downstream Evaluation
4.3 Quantitative Evaluation of Equivariance
In order to quantitatively evaluate the equivariant performance of our method and capsule projector, we employ commonly used metrics including Mean Reciprocal Rank (MRR) and Hit Rate at k (H@k). We utilise the same setup as described in [14], and discuss in further details the setup in appendix B.3. All the results evaluated by the aforementioned metrics are given in Table 3. Our CapsIE network outperforms EquiMod, Only Equivariance and SIE by a considerable margin across all metrics and for all dataset splits. We achieve strong perfromance on PRE, reporting 0.21 PRE on the validation set compared to 0.48 for EquiMod and Only Equivariance and 0.29 for SIE. The same large gains in equivariant performance are shown for MRR and H@1 and H@5. Note, a random H@1 results in a performance for 2% (0.02) demonstrating that our method lies well above random. Given that the same predictor is used for all reported methods, we can thus conclude how impactful the CapNet projector is at learning equivariant embeddings as opposed to linear projectors.
Table 3: Quantitative evaluation of the predictor when using a Capsule network projector, using PRE, MRR and H@k. The source dataset which embeddings are computed and the dataset used for retrieval are given in the format source-retrieval for PRE and source for MRR and H@k.
PRE ($\downarrow$) MRR ($\uparrow$) H@1 ($\uparrow$) H@5 ($\uparrow$)
Method train-train val-val val-all train val train val train val
EquiMod 0.47 0.48 0.48 0.17 0.16 0.06 0.05 0.24 0.22
Only Equivariance 0.47 0.48 0.48 0.17 0.17 0.06 0.05 0.24 0.22
SIE 0.26 0.29 0.27 0.51 0.41 0.41 0.30 0.60 0.51
CapsIE (ours) 0.17 0.21 0.20 0.60 0.47 0.50 0.36 0.71 0.58
@@ 7 Conclusion
Our proposed method demonstrates how CapsNets can be employed in self-supervised learning to better learn equivariant properties without the need to explicitly split representation vectors and train separate projector networks. The resulting solution, CapsIE, achieves state-of-the-art performance in equivariant and invariant downstream benchmarks with a significant improvement of 0.05 $R^{2}$ on prior self-supervised rotation prediction tasks and 0.02 $R^{2}$ improvement over the supervised baseline. In addition, we observe competitive performance of CapsIE on equivariant tasks not explicitly trained for, further demonstrating the implicit equivariant properties of our capsule architecture under standard invariant optimisation criteria akin to those of VICReg [3], SimCLR [7], and MSN [1]. Our results contribute significantly to the avocation of CapsNets in self-supervised representation learning, introducing desirable properties with improved effectiveness over MLP projectors.
@@ B.2 Downstream Evaluation
B.3 Quantitative Equivariant Evaluation Implementation Details
We evaluate the equivariant properties of the predictor in line with that proposed in [14] reporting the Mean Reciprocal Rank (MRR) and Hit Rate at k (H@k) on the multi-object setting. Given a source and target pose of an object, we first compute the embeddings of each image, and pass the source embedding through the predictor and use the resulting vector to retrieve the nearest neighbours.
The MRR is the average reciprocal rank of the target embedding in the retrieved nearest neighbour graph. H@k in this case is computed to be 1 if the target embedding is in the k-NN graph of the predicted embedding, where we only look for nearest neighbours among the views of the same object.
The Prediction Retrieval Error (PRE) gives an evaluation of predictor quality, and is given by the distance between its rotation $q_{1}\in\mathbb{H}$ and the target rotation $q_{2}$ as $d=1-<q_{1},q_{2}>^{2}$ of the nearest neighbour of the predicted embedding averaged over the whole dataset. Full implementation details can be found in the open-source code provided.
B.4 Supervised Training of SR-Caps